OpenStudy (anonymous):

If given f(x) = x + 4 and g(x) = –2x2 + 3x – 1, what is (f – g)(x)?

OpenStudy (anonymous):

(f-g)(x)= (f(x)-g(x))=[x+4]-[-2x2+3x-1]<--distribute the negative sign ===> x+4+2x2-3x+1(I believe you meant 2x^2) ===> 2x^2-2x+5